1.     The rate of heat transfer per unit area from a metal plate is 800 W. The surface temperature of the plate is 200°F, and ambient temperature is 68°F. Estimate the convective heat transfer coefficient.


1
Expert's answer
2021-05-10T08:27:22-0400

We can find the convective heat transfer coefficient from the formula:


"q=h\Delta T,""h=\dfrac{q}{\Delta T}=\dfrac{800\ \dfrac{W}{m^2}}{200\ ^{\circ}\!F-68\ ^{\circ}\!F}=6.1\ \dfrac{W}{m^{2}\cdot \!^{\circ}\!F}."